Lettuce Serve manufactures a variety of specialty salad dressings. Production runs are both high-volume and low-volume activities, depending on customer orders. Presented is Lettuce Serve's 2017 general manufacturing costs (manufacturing overhead) and each cost's related activity cost driver.

Level Total Cost   Units of Cost Driver
Unit $ 500,000   10,000 machine hours
Batch 50,000   500 customer orders
Product 100,000   25 products


Their cranberry vinaigrette dressing required 3,000 machine hours to fill 10 customer orders for a total of 8,000 units.

(a) Assuming all manufacturing overhead is estimated and predicted on the basis of machine hours, determine the predicted total overhead costs to produce the 8,000 units of cranberry vinaigrette.
$Answer


(b) Assuming manufacturing overhead is estimated and predicted using separate rates for machine hours, customer orders, and products (a multiple-level cost hierarchy), determine the predicted total overhead costs to produce the 8,000 units of cranberry vinaigrette.
$Answer

(c) Calculate the error in predicting manufacturing overhead using machine hours versus using multiple cost drivers. Indicate whether the use of only machine hours results in overpredicting or underpredicting the costs to produce 8,000 units of cranberry vinaigrette.
$Answer


(d) Determine the error in the prediction of cranberry vinaigrette batch-level costs resulting from the use of only machine hours. Indicate whether the use of only machine hours results in overpredicting or underpredicting the batch-level costs of cranberry vinaigrette.
$Answer
(e) Determine the error in the prediction of cranberry vinaigrette product-level costs resulting from the use of only machine hours. Indicate whether the use of only machine hours results in overpredicting or underpredicting the product-level costs of cranberry vinaigrette.
$Answer
Answer #1
WORKING OF ACTIVITY RATES:
Level Total Cost Units of Cost Driver Activity rate
Unit $500,000 10,000 machine hours $50 per MH
Batch 50,000 500 customer orders $                 100 per order
Product 100,000 25 products $             4,000 Per product
$650,000
OH assingment rate based on machine hour = 650000/10000 = $65 per MH.
a) Predicted total overhead costs on the basis of machine hours = 3000*$65 = 195000
b) Predicted overhead costs under multple-level cost hierarchy:
Machine operation = 3000*$50 = 150000
Customer order = 10*$100 = 1000
Product cost = 1*$4000 = 4000
Predicted overhead costs 155000
c) Error in prediction = 195000-155000 = 40000 Over prediction
d) Batch level costs using only machine hours = (50000/10000)*3000 = 15000
Batch level costs using multiple cost drivers 1000
Error in prediction = 15000-1000 = 14000 Over prediction
e) Product level costs using only machine hours = (100000/10000)*3000 = 30000
Product level costs using multiple cost drivers = 4000
Error in prediction = 30000-4000 = 26000 Over prediction